OpenStudy (anonymous):

g(x) = -x2 - 2x + 3 find the vertex.

OpenStudy (anonymous):

find the derivative,equate to 0 and find the values of x . I guess that will help .

OpenStudy (cwrw238):

if you havent done any calculus then do it this way x cood of vertex = b/2a = -2 / 2 = -1 y cood = -(-1)^2 - 2(-1) + 3 = 4 vertex is at (-1,4)